Find the x- and y-intercepts for the equation. Please show all of your work. Then graph the equation.
9y-3x=-6
I have one answer but don't know how to get more!
answer I have is (5,1)

Answer by mananth(16946)

9y-3x=6
Y intercept
plug x=0
9y=6
y=6/9=>2/3 (0,2/3)
X intercept
Plug y=0
-3x=6
x=-1 (-2,0)
plot the points and draw the line